A farmer has 3000 feet of wire to enclose a rectangular field. He plans to fence the entire area and then subdivide it by running a perpendicular fence across the middle. Find the dimensions of the field that would enclose the maximum area . What is the maximum area?
Can you please help me out? Thanks so much in advance:)

Answer by TimothyLamb(4379)

A = Lw
P = 2L + 2w = 3000
2L + 2w = 3000
2L = 3000 - 2w
L = 1500 - w
---
A = Lw
A = (1500 - w)w
---
A(w) = -w^2 + 1500w + 0
---
the above quadratic function is represented by the quadratic equation in standard form, with a=-1, b=1500, and c=0
---
to find the maximum area A(w), plug this:
-1 1500 0
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
Answer 1:
the maximum point of the above quadratic equation is: ( w= 750, A(w)= 562500 )
---
so the maximum area that can be enclosed is: 562500 sq.ft
---
now, to find the L and w dimensions of the maximum area, remember this equation from above:
A(w) = -w^2 + 1500w + 0
---
we used that equation to maximize Area, but as a side effect we also got w for the max area, from the max point of the quadratic equation ( w= 750, A(w)= 562500 )
---
Answer 2:
w for maximum area = 750
L for maximum area = 1500 - w = 1500 - 750 = 750
so the maximum area has dims: w=750, L=750 feet
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php